# Exploring Self-supervised Logic-enhanced Training for Large Language Models

### v4

@@ Abstract
[…]
@@ 5 Results and Analysis
Table 1: The results on logical reasoning benchmarks. Better results are annotated in bold. ${}^{{\dagger}}$ refers that the corresponding model is trained through QLoRA (Dettmers et al., 2023).
@@ 5.1 Logical Reasoning
    w/ ${}^{{\dagger}}$ 68.8 68.1 54.4 58.3
Table 2: The results of LLaMA models on RACE and MMLU. ${}^{\dagger}$ means training through QLoRA.
@@ 5.2 Hybrid Reasoning and Application
    w/ ${}^{{\dagger}}$ (no aug.) 49.4 53.0 44.2 40.8
    w/ ${}^{{\dagger}}$ 50.2 54.4 45.9 42.6
Table 3: The effect of different training objectives. Ctr refers contrastive learning and ar means the auto-regressive variant. no aug. means the counterfactual data augmentation is removed from the LogicLLM framework. ${}^{{\dagger}}$ means that the model is trained with QLoRA.
@@ 5.3 Pre-training Strategy
Moreover, the auto-regressive objective significantly reduces computation costs during training by eliminating the need for negative candidates encoding. The streamlining of training process leads to more efficient and resource-friendly training without sacrificing performance.
@@ 5.6 Data Assumption Auto-Verification
Model Normal Normal (Anony.) Count. Count. (Anony.)
@@
Table 5: The ratio of consistent data deemed by ChatGPT and GPT-4. Anony. refers to anonymization and Count. is the simplification of Counterfactual.
@@ Appendix B Interpretation for Different Results on RACE
Appendix C More Experimental Results
C.1 Logic-enhanced Meta-training for Complex Task Understanding

### v6

@@ Abstract
Fangkai Jiao1,2   Zhiyang Teng1   Bosheng Ding1   Zhengyuan Liu2 Nancy F. Chen2, [^2] Correspondence to: Nancy F. Chen and Shafiq Joty.   Shafiq Joty1,3, [^2] Correspondence to: Nancy F. Chen and Shafiq Joty. 1Nanyang Technological University, Singapore 2Institute for Infocomm Research (I2R), ASTAR, Singapore  3Salesforce Research jiaofangkai@hotmail.com  bosheng001@e.ntu.edu.sg  zhiyang.teng@ntu.edu.sg {nfychen, liu_zhengyuan}@i2r.a-star.edu.sgsjoty@salesforce.com
@@ 5 Results and Analysis
Table 1: The results on logical reasoning benchmarks. Better results are annotated in bold. refers that the corresponding model is trained through QLoRA (Dettmers et al., 2023).
@@ 5.1 Logical Reasoning
       w/ LogicLLM 68.8 68.1 54.4 58.3
Table 2: The results of LLaMA models on RACE and MMLU. means training through QLoRA.
@@ 5.2 Hybrid Reasoning and Application
       w/ LogicLLM (no aug.) 49.4 53.0 44.2 40.8
       w/ LogicLLM (1 aug.) 50.8 52.7 45.6 41.5
       w/ LogicLLM 50.2 54.4 45.9 42.6
Table 3: The effect of different training objectives. Ctr refers contrastive learning and ar means the auto-regressive variant. no aug. means the counterfactual data augmentation is removed from the LogicLLM framework.  means that the model is trained with QLoRA.
@@ 5.3 Pre-training Strategy
Moreover, the auto-regressive objective significantly reduces computation costs during training by eliminating the need for negative candidates encoding. The streamlining of training process leads to more efficient and resource-friendly training without sacrificing performance. We also add another experiment by adjusting the ratio between counterfactual data and the normal ones as 1:1, and the comparison reveal that mixing more counterfactual data can also benefit the performance, which could be especially useful for low-resource domain, like finance and multi-lingual LLMs.
@@ 5.6 Data Assumption Auto-Verification
Model Normal Normal (Anony.) C.F. C.F. (Anony.)
@@
Table 5: The ratio of consistent data deemed by ChatGPT and GPT-4. Anony. refers to anonymization and C.F. is the simplification of Counterfactual.
@@ 6 Conclusion
Acknowledgements
This research is supported by the Ministry of Education, Singapore, under its Science of Learning Grant (award ID: MOE-MOESOL2021-0006). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of the Ministry of Education, Singapore.
Besides, we sincerely appreciate the valuable comments from all the reviewers to help us make the paper polished. We also greatly thank to Chengwei Qin and Professor Aixin Sun for their kind suggestions.
@@ Appendix B Interpretation for Different Results on RACE
Appendix C Logic-enhanced Meta-training for Complex Task Understanding